a beam of light converges toa point S. a mirror is placed in front of the screen at a distance of 10cm from the screen. it is found that the beam now converges at a point 20cm in front of the mirror. find the focal lenght.
i did not receive any answer to this question.

Grade:10

1 Answers

Vikas TU
6 years ago
Dear Student,
Given data
Without mirror beam of light converges at point S.
When the mirror is placed in front of mirror at point O then OS = 10 cm
Then the object distance that is u = 10 cm
And now it converges at point 20 cm that is  v = 20 cm
According to the mirror formula:
1/v +1/u = 1/f
1/20+1/10 = 1/f
3/20 = 1/f
F = 20/3
Focal length = 6.67 cm
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)